A uniform electric field with 500 N/C is parallel to a flat square area of 115 m 2 . Calculate the electric flux.


1
Expert's answer
2022-02-08T13:03:27-0500

Given:

"E=\rm 500 \: N/C"

"A=115\:\rm m^2"

"\theta=90^\circ"


The electric flux through the surface

"\phi=EA\cos\theta=500*115*\cos90^\circ=0"